OpenStudy (anonymous):

simplify x^2*x^3, by first expanding X^2 and x^3

OpenStudy (turingtest):

(x*x)(x*x*x)=x*x*x*x*x=x^5

OpenStudy (saifoo.khan):

\[\color\red{x \times x} \times \color\green{x \times x \times x }\]

OpenStudy (saifoo.khan):

\[x^{1 + 1 + 1+ 1+1}= \Huge x^5\]